If A and B are independent events, P(A) = 0.3, and P(B) = 0.7, determine P(A∪B).

A. 0.21

B. 0.40

C. 0.79

D. 1.00
Answer #1

P(A) = 0.3

P(B) = 0.7

For independent events A and B, P(A \cap B) = P(A) x P(B)

= 0.3 x 0.7

= 0.21

P(A U B) = P(A) + P(B) - P(A \cap B)

= 0.3 + 0.7 - 0.21

= 0.79

Ans: c. 0.79